OpenStudy (abmon98):

solve the equation 3cos2x-4sin2x=2 0

ganeshie8 (ganeshie8):

there is a direct formula for representing `acosx + bsinx` as a single sinusoidal function

OpenStudy (abmon98):

\[3(\cos^2x-\sin^2x)+4(2sinxcosx)=2\] \[8sinxcosx+3-6\sin^2x=2\]

ganeshie8 (ganeshie8):

say 2x = t \[3\cos t - 4 \sin t = \langle 3, -4\rangle \bullet \langle \cos t, \sin t \rangle \]

OpenStudy (abmon98):

but isnt that a double angle?

ganeshie8 (ganeshie8):

lets hope the substitution works :) if it doesn't work, we can try the other method

ganeshie8 (ganeshie8):

\[= |\langle 3, -4\rangle | ~ | \langle \cos t, \sin t \rangle | \cos (t - \phi) \] \[= 5 \cos (t - \phi)\] \(\phi = \tan^{-1}(\frac{-4}{3})\)

OpenStudy (abmon98):

Cos-1=2/5=66.42182 Cosine is positive in the 4th quadrant 360-66.42182=293.578 293.578+tan-1(-4/3)=346.708

OpenStudy (abmon98):

66.42182+53.130=119.55

OpenStudy (abmon98):

although iam not sure if my working is right or not?@ganeshie8

OpenStudy (abmon98):

@ganeshie8 Could you help in solving with the other method if you are free?

ganeshie8 (ganeshie8):

solve : \[5\cos\left(2x -\tan^{-1}(\frac{-4}{3})\right) = 2\]

ganeshie8 (ganeshie8):

\[2x -\tan^{-1}(\frac{-4}{3}) = \cos^{-1}(\frac{2}{5})\]

ganeshie8 (ganeshie8):

\[2x -(-53.13^{\circ}) = 66.42^{\circ}\]

ganeshie8 (ganeshie8):

\[2x = 66.42^{\circ} - 53.13^{\circ}\]

ganeshie8 (ganeshie8):

\[2x = 13.29^{\circ}\]

ganeshie8 (ganeshie8):

since cos(x) = cos( \(\color{red}{-}\)x), the general solutions are : \[2x = 13.29^{\circ} + 360 n\] and \[2x = \color{red}{-}13.29^{\circ} + 360 n\]

ganeshie8 (ganeshie8):

plugin n = 0,1,2,... and solve x

OpenStudy (abmon98):

why is it -13.29

ganeshie8 (ganeshie8):

the short answer is : cos(x) = cos(-x)

ganeshie8 (ganeshie8):

there fore, below two equations are one and same : cos(x) = 2 cos(-x) = 2

ganeshie8 (ganeshie8):

you may verify your answer with wolfram : http://www.wolframalpha.com/input/?i=solve+3cos%282x%29-4sin%282x%29%3D2%2C++0%3Cx%3C2pi

ganeshie8 (ganeshie8):

you will need to covnert the solutions to degrees ^

OpenStudy (abmon98):

Thank you @ganeshie8

ganeshie8 (ganeshie8):

np :)